OpenStudy (anonymous):

I give medals! i need to prove this identity:

OpenStudy (anonymous):

\[\cos (\frac{ \pi }{ 4 } +x)+\cos (\frac{ \pi }{ 4 }-x)= \sqrt{2} cosx \]

OpenStudy (anonymous):

@jhannybean

OpenStudy (jhannybean):

Hm, there's an identity that falls with cosines and the addition or subtraction of angles. do you know what it is?

OpenStudy (anonymous):

yes one second

OpenStudy (anonymous):

cos(A+B) = (cosA)(cosB) + (sinA)(sinB) with this formula, it could be + or - based on the way it needs to be used

OpenStudy (jhannybean):

Yes :)

OpenStudy (jhannybean):

So we could use the cosine subtraction rule first \[\cos(A-B) = \cos(A)\cos(B)+\sin(A)\sin(B)\]\[\rm cos\left(\pi/4 -x\right) = \cos(\pi/4)\cos(x) +\sin(\pi/4)\sin(x)\]This completes the first part of the identity.\[\boxed{\cos(\pi/4)\cos(x) + \sin(\pi/4)\sin(x)} +\cos\left(\pi/4) + x\right) = \sqrt{2}\cos(x)\]

OpenStudy (anonymous):

okay so far so good

OpenStudy (jhannybean):

Now evaluate \(\cos\left(\frac{\pi}{4}\right)\) and \(\sin\left(\frac{\pi}{4}\right)\)

OpenStudy (anonymous):

and that is the \[\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (jhannybean):

Or otherwise written as \(\frac{1}{\sqrt{2}}\)

OpenStudy (anonymous):

so i now need to make that 1 disappear right?

OpenStudy (jhannybean):

So from this step\[\rightarrow \boxed{\cos(\pi/4)\cos(x) + \sin(\pi/4)\sin(x)} +\cos\left(\pi/4) + x\right) = \sqrt{2}\cos(x)\] we can substitute the values we've just found and simplify. \[\frac{\cos(x)}{\sqrt{2}}+\frac{\sin(x)}{\sqrt{2}} +\cos(\pi/4 +x) =\sqrt{2}\cos(x)\]

OpenStudy (jhannybean):

and like you mentioned earlier, we have our identity for the addition of angles within the cosine function. \[\cos(A+B) =\cos(A)\cos(B) -\sin(A)\sin(B)\]\[\cos(\pi/4 + x) = \cos(\pi/4)\cos(x)-\sin(\pi/4)\sin(x)\]

OpenStudy (anonymous):

okay, im just reading through what youve put up. its starting to piece together and make more sense

OpenStudy (jhannybean):

So you end up with the same result. \[\frac{\cos(x)}{\sqrt{2}}+\frac{\sin(x)}{\sqrt{2}}+\boxed{\cos(\pi/4)\cos(x)-\sin(\pi/4)\sin(x)} =\sqrt{2}\cos(x)\]

OpenStudy (anonymous):

my original problem only involved cos. Since sin is in the answer, is that a valid answer?

OpenStudy (anonymous):

also, basically we started with only cos, but in order to get to the answer, we had to involve sin

OpenStudy (anonymous):

is that right?

OpenStudy (jhannybean):

Yes, the sine function only came into the picture because of the cosine addition and subtraction rules.

OpenStudy (anonymous):

ohh okay, thank you :)

OpenStudy (jhannybean):

\[\rightarrow \frac{\cos(x)}{\sqrt{2}}+\frac{\sin(x)}{\sqrt{2}}+\boxed{\cos(\pi/4)\cos(x)-\sin(\pi/4)\sin(x)} =\sqrt{2}\cos(x)\]\[\frac{\cos(x)}{\sqrt{2}} +\frac{\sin(x)}{\sqrt{2}}+\frac{\cos(x)}{\sqrt{2}}-\frac{\sin(x)}{\sqrt{2}}=\sqrt{2}\cos(x)\]

OpenStudy (jhannybean):

You see how the sine functions cancel eachother out?

OpenStudy (anonymous):

yes, because of their signs (- and +), right?

OpenStudy (jhannybean):

Yes :)

OpenStudy (jhannybean):

Now we're left with \[\frac{2\cos(x)}{\sqrt{2}}=\sqrt{2}\cos(x)\]and we know \(2 =\sqrt{2}\cdot \sqrt{2}\) so the fraction disappears, and we're left with\[\sqrt{2}\cos(x) = \sqrt{2}\cos(x)\]:D!

OpenStudy (anonymous):

i love you. thank you. so much. could you double check just one more, i promise? i already solved it, can i check with you (under a different post) to confirm that i did it correctly? youre extremely smart. ill give you another medal (a third medal is pretty much the equivalent of a trophy ;) )

OpenStudy (anonymous):

wow that really clarified everything

OpenStudy (jhannybean):

haha thank you! Sure, i you liek, rework the problem and I shall point out your mistakes. Keep this tab open so you can cross reference my work to yours :P

OpenStudy (anonymous):

okay, you sure? i just want to credit you where due

OpenStudy (jhannybean):

I'm sure :)

OpenStudy (anonymous):

this is difficult to put on here, im going to upload the math i did for it lol

OpenStudy (jhannybean):

Alllrighty :) Take a picture of your work on paper! :)

OpenStudy (anonymous):

i am :) just waiting for it to download to my slow computer -_-

OpenStudy (anonymous):

okay almost don. sorry it wsant sending to my email. 2 seconds

OpenStudy (anonymous):

sorry its illegible, can you make out the problem?

OpenStudy (jhannybean):

Think you did something wrong there, it should come out to \[\boxed{\frac{\sqrt{3}}{2}}\]

OpenStudy (anonymous):

Yeah, I felt like I did. Do you mind explaining how you got that?

OpenStudy (jhannybean):

Working on it, burning myself out a bit, lol.

OpenStudy (jhannybean):

Okay, it looks like you were making it a little harder than it actually was.

OpenStudy (anonymous):

Okay, no problem, sorry. Btw if it looks like I sign out and sign back in or something its because my computers being a jerk at the moment.

OpenStudy (anonymous):

That should be my middle name when it comes to math actually. I always make it harder than it is lol

OpenStudy (jhannybean):

\[\cos^2(15) - \sin^2(15)\] This identity states \[\cos^2(x)-\sin^2(x) \implies 2\cos^2(x) -1 \implies 1-2\sin^2(x) \implies \cos(2x)\]

OpenStudy (jhannybean):

So there are various forms you can use in solving for whatever you're solving for.

OpenStudy (jhannybean):

I personally like using \(2\sin^2(x)-1\) when I'm given a problem that relates to the double angle formula involving cosine, so I used that.

OpenStudy (anonymous):

You've been insanely helpful this entire time, so I'm with you on that version. it's a-ok by me.

OpenStudy (jhannybean):

\[\cos^2(15) -\sin^2(15)\] Let's change the degree to radians. You do this by :\[15~ \text{deg} \cdot \frac{\pi}{180 ~\text{deg}} = \frac{\pi}{12} \]

OpenStudy (x3_drummerchick):

wow this guy is smart

OpenStudy (nikato):

Sorry to interrupt... But I found an even shorter method Using double argument cos^2 (x) - sin^2 (x) =cos(2x) And from your problem, you see that x=15 So all you have to find is cos(2x15) =cos(30) And using the unit circle or whatever, You know cos(30)= sqrt3/2

OpenStudy (jhannybean):

Yeah that works too :)

OpenStudy (anonymous):

Hmm, okay, thanks yeah as long as I understand the method behind the madness

OpenStudy (anonymous):

I was gonna say jhannybean, you've stuck by me on this stuff, and I understand the way you were explaining it, so you can complete it your way if you want as well

OpenStudy (jhannybean):

So we have : \[\cos^2\left(\frac{\pi}{12}\right) -\sin^2\left(\frac{\pi}{12}\right) \]\[=2\cos^2\left(\pi/12\right) -1\]\[=\frac{\sqrt{3}}{2}\]

OpenStudy (jhannybean):

I'm sorry I cannot help you with any more problems after this, I'm totally burnt out and a bit exhausted.

OpenStudy (anonymous):

I totally understand, I appreciate all you've done, thanks

OpenStudy (jhannybean):

What I suggest you do is take a gigantic index card and write out all the formulas you are learning in class. When you do your homework, instead of looking through your book for the identities, you will have your flashcards with you.

OpenStudy (jhannybean):

That way not only will you lern the formulas, you can memorize and apply them later on with ease. Atleast that's how I went about doing my Trigonometry homework. :)

OpenStudy (jhannybean):

learn*

OpenStudy (anonymous):

That makes sense, I'll def do it! Thanks a million. A thousand gold trophies for you!!!! THOUSANDS!

OpenStudy (jhannybean):

Haha thanks. Good luck with learning the rest of your identities! :-)

OpenStudy (anonymous):

Thank you!! Take care!